OpenStudy (z4k4r1y4):

A square loop of wire is held in a uniform 0.34 T magnetic field directed perpendicular to the plane of the loop. The length of each side of the square is decreasing at a constant rate of 4.9 cm/s. What emf is induced in the loop when the length is 15 cm?

OpenStudy (z4k4r1y4):

I got: \[\frac{ dA }{ dt }=2.4mm ^{2}\] Then I don't know where to go from there.

OpenStudy (z4k4r1y4):

how do I get to Faraday's Law from there?

OpenStudy (michele_laino):

Let's suppose that the initial length of each side of the square is \(l_0\), then we can write: \[l\left( t \right) = {l_0} - vt\] where \(v=4.9\), and the corresponding area \(S(t)\), as a function of time \(t\), is: \[S\left( t \right) = {\left( {{l_0} - vt} \right)^2}\] Next we can write, the magnetic flux \( \Phi(t)\) of field \(B\), is: \[\Phi \left( t \right) = BS = B{\left( {{l_0} - vt} \right)^2}\] and the corresponding induced voltage \(E\) ( \(Farady\) law,s), is: \[E = - \frac{{d\Phi }}{{dt}} = 2Bv\left( {{l_0} - vt} \right)\] of course at time \(\tau\) the length is \(l_1=l_0-v \tau=15\), so we can write: \[E\left( \tau \right) = 2Bv\left( {{l_0} - v\tau } \right) = 2Bv{l_1} = ...?\] please continue

OpenStudy (z4k4r1y4):

Thank you very much. I was struggling to derive an equation for the magnetic field with respect to time.